Finance Functions If I invest $2,100 a year for 40 years toward my retirement and earn 8 percent a year on my investments, hohow do I do this in excel
Answer #1

1. Making deposits at the beginning of the year:

=FV(rate,nper,pmt,pv,0 or 1)

=FV(8%,40,-2100,,1)

=587,540.18

2. Making deposits at the end of the year:

=FV(8%,40,-2100,,0)

=544,018.69

To build a model, you can assign different cells for different parameters as in rate A1, nper A2 etc. and the function will look like =FV(A1,A2,A3,,1)